OpenStudy (anonymous):

Find the equation for the normal line to f at the point (-4, f(-4)). f(x)=-3-2x^3

OpenStudy (anonymous):

didn't we just do this one?

OpenStudy (anonymous):

slightly diff

OpenStudy (amistre64):

more elevation on the question tier :)

OpenStudy (anonymous):

I got y+125=(-1/96)(x+4)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

once you guys explain, I have to check if I understand by doing another example

OpenStudy (anonymous):

getting it correct is one thing, understanding another

OpenStudy (amistre64):

f(-4)=-3-2(-4)^3 f'(x)=-6x^2; f'(-4) = -6(-4)^2 solve these and your good to go really

OpenStudy (anonymous):

steps exactly the same as before. find \[f(-4)\] for the second coordinate \[f'(-4)\] for the slope and then take the negative reciprocal for the slope of the perpendicular line etc

OpenStudy (amistre64):

f(-4) = 125; f'(-4) = -96 right?

OpenStudy (amistre64):

y = -(1/-96)(x+4)+125 is how to fill in the equation, then simplify as need be

OpenStudy (anonymous):

yes -96

OpenStudy (anonymous):

and 125

OpenStudy (anonymous):

does that mean I was wrong?

OpenStudy (amistre64):

equation of the normal at x=-4 is: y = (1/96)x +3001/24 if I did it right

OpenStudy (anonymous):

can you do it in the other form too pls: is y +125 or y - 125? etc

OpenStudy (amistre64):

y - 125 = m(x+4) ; where m = (1/96)

OpenStudy (anonymous):

i sm not minus?